# EvoC2Rust: A Skeleton-guided Framework for Project-Level C-to-Rust Translation

### v2

@@ 6.2. Project-level Code Translation with LLMs
Shiraishi et al. (Shiraishi and Shinagawa, 2024) pioneered this direction by using project metadata to maintain consistency, though their method prioritizes compilation success over functional correctness. Syzygy (Shetty et al., 2024) and RustMap (Cai et al., 2025) employ a strategy of translating both functions and test cases while utilizing feedback mechanisms for repair. Nevertheless, these approaches necessitate complete dependency contexts (occasionally through manual intervention) and exhibit limited scalability when applied to complex industrial codebases. Khatry et al. (Khatry et al., 2025) and Ou et al. (Ou et al., 2024) introduced repository-scale benchmarks and self-repair techniques, yet rely on annotated Rust interfaces rarely available in practice. Hong et al. (Hong and Ryu, 2025b) addressed this with Tymcrat, a type inference-based method for automating signature translation.

### v3

@@ 6.2. Project-level Code Translation with LLMs
Shiraishi et al. (projectc2rust) pioneered this direction by using project metadata to maintain consistency, though their method prioritizes compilation success over functional correctness. Syzygy (Syzygy) and RustMap (RustMap) employ a strategy of translating both functions and test cases while utilizing feedback mechanisms for repair. Nevertheless, these approaches necessitate complete dependency contexts (occasionally through manual intervention) and exhibit limited scalability when applied to complex industrial codebases. Khatry et al. (CRUST-Bench) and Ou et al. (RustRepoTrans) introduced repository-scale benchmarks and self-repair techniques, yet rely on annotated Rust interfaces rarely available in practice. Hong et al. (typemigrating) addressed this with Tymcrat, a type inference-based method for automating signature translation.
@@ 7. Conclusion
Acknowledgment
This research is funded by the National Key Research and Development Program of China (Grant No. 2023YFB4503802), the National Natural Science Foundation of China (Grant No. 62232003), and the Natural Science Foundation of Shanghai (Grant No. 25ZR1401175).